OpenStudy (anonymous):

Sean earns $4.50 per hour more than his trainee, Tyrone. During one 8 hour day, they earn a total of $284. How much does Tyrone make per hour?

OpenStudy (btaylor):

s = hours Sean works t = hours Tyrone woks s = t + 4.50 8s + 8t = 284 8(t+4.50)+8t=284 16t+36=284 16t=248 t = $15.50

OpenStudy (anonymous):

You are really smart. Just letting you know:p